OpenStudy (anonymous):

how do i find the geometric mean between 4 and 25

hartnn (hartnn):

if b is the geometric mean between a nd c, then \(\large b=\sqrt{ac}\) here, a=4,c=25

OpenStudy (anonymous):

multiply (in your head) then take the square root (again in your head)

OpenStudy (anonymous):

can you show me how i would set it up as 2 fractions?

hartnn (hartnn):

as a fraction ? yes, if b is the geometric mean between a and c, then \(\huge \dfrac{a}{b}=\dfrac{b}{c}\)

OpenStudy (anonymous):

a and c is 4 and 25?

hartnn (hartnn):

absolutely :)

OpenStudy (anonymous):

so i would set it up as 4/b=b/25 ?

hartnn (hartnn):

correct, go on

OpenStudy (anonymous):

and the sqrt of 4*25 is 10?

hartnn (hartnn):

yes, it is. GM =10 is correct :) any doubts ?

OpenStudy (anonymous):

nope :) but what if instead the 25 was 1/4 as a fraction?

hartnn (hartnn):

but what if ..... what ? i couldn't get you....

OpenStudy (anonymous):

this a new question lol

hartnn (hartnn):

ohh...i'll go there.. :)